if the area of a circular pond is 153.86 sq. yd. what is the circumference of the pond?
Answer by stanbon(75887)

if the area of a circular pond is 153.86 sq. yd. what is the circumference of the pond?
-------------------------
Area = (pi)r^2 = 153.86 sq. yd.
r^2 = 153.86/pi
r = sqrt(48.98)
-----
Circumference = 2(pi)r
C = 2(pi)sqrt(48.98)
C = 43.97...
Cheers,
Stan H.